The water in the pipes of a typical house is at a gauge pressure of 2.9 times atmospheric pressure. What is the absolute pressure in pascals and in units of pounds per square inch? Pa xlb/in2
Answer #1

Value of atmospheric pressure, P0 = 101325 Pa

Pressure of water in the pipe, P1 = 2.9 x P0 = 2.9 x 101325 Pa = 293842.5 Pa

Gauge pressure attached to the water pipes measure pressure with respect to the atmospheric pressure.

Definition of absolute pressure of a point is equal to the gauge pressure and atmospheric pressure.

So, absolute pressure of water pipe = P1 + P0

= 293842.5 + 101325 = 395167.5 Pa (Answer)

Now, convert this pressure into pound per square inch -

Absolute pressure, P = 395167.5 x 0.00015 = 59.27 lb/in^2 (Answer)